40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11224     Entry time: Thu Apr 16 10:32:42 2015
Author: Steve 
Type: Update 
Category: Optical Levers 
Subject: oplevs monitoring 

The IFO_overview of oplevs seems ok, The servos are working fine. The green arms are locked, but master and oplev_summary monitoring screens are not.

I'm proposing to Erick G. to postpone the oplev noise measurement.

Attachment 1: oplevCond.png  129 kB  | Show | Show all
ELOG V3.1.3-